VIS Meeting on 2016/5/13
2016/5/13 13:30 -
Participants:Hirata, Takahashi, Ishizaki, Mark, Fujii, Okutomi, Fabian, Aso, Tanaka, Sato, Shoda, Akutsu
The current overall schedule chart.
iKAGRA related activities
PR3 post-run work status & plan (Shoda)
- Transfer function measurements in the air: same as before.
0.2Hz peak in OpLev was caused by the increased seismic after the earthquake.
- Locked the suspension to measure the 3D coordinates of the mirror
- 5/16: measurement of PR3 mirror position relative to the vacuum tank
5/17: OpLev wind shield test?, Inner frame status check.
- 5/18: Extended chief meeting
- 5/19, 20: Evacuation
IP tests in Kamioka
- Status report by Fujii: slides
- Control test may be performed with the hanging test of BS
- Ask if we can perform the control test remotely (Aso)
- Ask Joris what is his intention and schedule (Raffaele)
iKAGRA ETM suspension
- Uninstallation status (Ohishi)
- Mirrors were shipped to Mitaka ?
Length sensing OpLev
- #In order to test this with the BS test haning, we want to have wide OSEMs as witness sensors on the mirror.
- #Stab-flag and tipped flag are both in quotation. We will order both. The gluing jigs will be the same.
Who will design and prepare the BS length sensing OpLev ? -> Ask Simon (Aso)
- What shall we do with PR2
PR2 chamber has only one viewport avaibale at 45deg. angle for OpLev
Original idea: Put a reflecting mirror on the bread board and monitor the rotation of the bread board with another OpLev.
- No length sensing
Two OpLevs needed
Another idea: Introduce the light with a fiber into the vacuum chamber to form a one-way 45deg. OpLev for length sensing.
- Length sensing possible
- Affected by the vacuum chamber tilt (if we mount the fiber collimator to the inner frame, the tilt will be less).
- Need to find a place to mount the fiber collimator inside the vacuum chamber
See if we can find any mounting point on the inner frame (Akutsu) by the next next VIS meeting.
- Suspended bread board
- #Many critical components will be installed on it (baffles, steering mirrors for the green injection etc)
- We have to make sure that the bread board suspension is properly damped
- Takahashi checked Sekiguchi thesis
Looks like we can estimate the damping force from his experimental results. Takahashi will calculate the necessary number of magnets.-> by the next VIS meeting
- Accelerometers for IP
- Geophone is too noisy for innertial damping in Kamioka
- TAMA SAS accelerometers for Type-A
- NIKHEF accelerometers for BS ?
- Which accelerometers for SR ?
- Buy from NIKHEF ? How much is the cost ?
- Get the design from NIKHEF and produce them in Japan (ATC) ?
- Pre-isolator assembly and tests
- What is the plan for SR Type-B
- Takahashi and Mark will start the discussion on, where, who, when.
- Bottom filter local sensor for Type-A
- What is the plan ?
- Tell SEO about the wider cross tube
- EQ stop design should take into account of the sensor/actuator mounts.
- Fix EQ stop at the bottom ?
- Use the same sensor/actuator as Type-Bp ?
- Sato-san is responsible for this part.
- What is the plan ?
- #Dummy payload design for Type-A
- Okutomi-kun is assigned.
- Wide OSEM production (Akutsu)
- For BS test hanging, we need the wider version by mid-to-late June.
- Got cost estimates for the small parts. Will order for 20 OSEMs.
- What are necessary in the next few weeks ?
- We need to count the number of necessary circuits for Type-A and Type-Bp, so that we can tell AEL the total number necessary for VIS.
Make a list by the end of May (Shoda, Takahashi)List
- There are some uncertainties in the number of necessary circuits.
- Make a list with ranges of numbers for uncertain things by the next VIS meeting. Then talk to AEL.
SRx will probably not have geophones, so Mark will need to do a revision too.
- Watch dog system (Tanaka)
- photo-MOS relay switch resistance fluctuation
Noise requirement for the resistance fluctuation is necessary (Aso)
Tanaka measured the photo-MOS relay resistance fluctuation-> 37mOhm RMS upper limit
- Will make spectrum of the resistance fluctuation.
DC-DC convertor noise check: quiter than a battery ?
#The next step: Make a prototype with slow shutdown mechanism
- Past Week Summary
- Flags and metal prisms ordered; expected 5/31.
New assembly drawing D1605179 focussing on items not part of main chain (and associated screws):
- Crane adapter
- Lower breadboard blade module (with cable clamps)
IM-RM safety bar (D1605127 from last week now obsolete)
- 3D assembly changed to have geophones instead of accelerometers (but SRx will probably be per original design - need to revisit circuit diagram)
Concept for load gauge adapter decided (reuse crane adapter plate and jacks from below PI; add Misumi crossbars, cleanroom-compatible chain hoist: Crane Frame.pdf)
Concept for second floor of clean booth decided (remove all floor panels; add bridge panels at right angles: BS Cleanbooth.pdf)
- Modified drawings for main assembly frame and drawings for most small parts (jack adapters etc) received from Mirapro.
- Wire sizes decided: 300 µm piano wire and 500 µm tungsten (will revisit this later).
- Hirata-san measured another SF.
- Need more realistic schedule
- Take into account the preparation of SR suspensions
- SRM and PRM payloads need to be ordered.
- IM assembly
- Planning to do it in ATC. Use the clean room used for the SF assembly.
- Fabian will start the preparation from the next week.
- Can we make SR wire breakers bigger?
Fukushima-san preparing test: 3/4/2016 + "1 month" = ≈4/1/2016, not ready as of 4/8.-> Takahashi will ask the status.
- Flags, Prisms
- BS dummy optic
Ordered, expected ≈4/22 (T1604811)
- Decided to have it shipped to Mitaka for initial assembly and give it to Tatsumi-san for prisms/flags.
- Mirror box (Tatsumi)
- Tatsumi-san has ordered modifications to mirror box for final flags and prisms.
- An extra structure attached to the EQ stop to prevent tipping over of the BS.
- Will consider if we can make it to the test hanging. But not critical for the test hanging.
- No action since last week.
- Not enough crane height problem (Hirata, Saito)
- No action since last week
- EQ stop screws, extensions and lock nuts (in manufacture - promised for 4/22)
- Ballast masses for payload
Hirata-san created drawings DropBox link
- Drawings were checked by Mark, not yet ordered.
- Lower breadboard stuff
- Breadboard spring bases and associated cable clamps (in manufacture, expected 4/15)
- Breadboard springs (in manufacture)
- Damper ring mods (in manufacture, expected 4/28)
- Nail-head parts: quote from Mirapro received, not urgent, will be combined with assembly frame order.
- BS Assembly Frame Design (Mark, Hirata)
- Hirata-san did new Inventor assembly with Mirapro's initial design plus various improvements including clean bag supports.
- Clean bag
- Ordered - expected 6/8.
- 2D and 3D CAD
- BS 3D Assembly
- 3D CAD for BS assembly frame revised again (seismometers changes to geophones).
- Item list and timetable was given to Miyakawa-san - some corrections received and added to schematic but not yet posted.
- Status on AEL boxes OK according to Miyakawa-san.
- Need to order LVDT oscillator cards from NIKHEF (Takahashi) and have a signal generator and cable in the meantime (Mark).
- Need to get new longer versions of some previously ordered BF-PI and SF-PI cables (Takahashi).
- Need to check previously ordered TM-BF and IM-BF OSEM cables have been flipped (Takahashi).
- Decided to hand-make cables for geophones because previously ordered cables have no conductor on pin 5.
Stepper drive box (requested by Sekiguchi-san T1503385) is being designed, should be ready in time.
- Procedure documents
Mark worked on BS Assembly Procedure E1504235 draft
Task list (Takahashi)
- Where to perform the TF+IP assembly ?
- The large access door to the ATC clean room is not available until Octorber or November. Doing the assembly work in Mitaka is next to impossible.
Prototype testof the EQ stop
- Purpose of the test needs to be clarified to decide on where to perform the test.
- The test of prototype long maraging rod was finished
- Will order them with maraging rods for the cryogenic payload early July. The lead time is 2 months.
- Type-B Standard Filters preparation
5 SF measurements were done. http://gwdoc.icrr.u-tokyo.ac.jp/cgi-bin/private/DocDB/ShowDocument?docid=5105
- The shell of the 2 filters were too tight. Were able to open but could not close one of them.
- The ones we cannot close the shell, we can use them for the Type-Bp.
- Ship them to Kamioka
- Type-A SF preparation
- Needs to cut more finshing rod blades.
- The next Type-A meeting will be on 5/26.
- #Simultion (Okutomi)
- #Local sensors for BF
- Modification for bKAGRA
- Discussing with ATC people for the mechanical modification.
- Got some comments
- Estimating the noise requirements for the BF sensor/actuator
- Full CAD model of the original Type-Bp needs to be created
Site work plan for the next week
Mon:PR3 measurements, IP
- Tue: PR3 measurements, IP
- Wed: Extended Chief Meeting
- Thu: PR3 measurements, IP
- Fri: PR3 vacuum evacuation, IP
- Travel (Week of 5/16)
- Takahashi: 17-20 (chief meeting and IP)
- Travel (Week of 5/23):
- Shoda: Elba
- Okutomi: Elba
Fujii:23-27 (Accelerometer, PR3)
Mark is away: July 8-20
- #GAS blade inspection by Virgo (Takahashi)
- Shipped the blades. Will get results in 1-2 months.
2016/5/20(Fri) 13:30 -